36-303: Sampling, Surveys & Society HW01: Due Tue Jan 25, 2010 in class

\underline{Reminders:}

• Things to do:

Also Due Tues jan 25: Team Assig. I.1 (two project proposals). See handouts “project-schedule.pdf”

and “project-ideas.pdf” on the class website, http://www.stat.cmu.edu/~brian/303.

Will be assigned in HW02: http://www.cmu.edu/osp/regulatory-compliance/human-subjects.html

lists CMU IRB application forms and various training materials for ethical conduct of human subjects research. Under “Education Requirements” there is a link to “CITI’s website”, the training website of the Collaborative Institutional Training Initiative (CITI).

1. Follow the “CITI’s website” link.

2. Select Carnegie Mellon University as your “participating institution” and create an account.

3. When you have created an account you will be directed to a page titled ”Select Curriculum” which displays a list of courses on Human Subjects Research: Biomedical Research Investigators; Social & Behavioral Investigators, and IRB Members. Choose the Social & Behavioral Investigators human subjects research module.

4. The course may take a few hours to complete but can be done over a period of time. When you complete the course, CITI will e-mail your completion record to the CMU IRB.

5. From the CITI main menu, you can also print a copy of the completion report for this course. Turn in a printed copy of your completion report with this homework.

If you already have an active certificate of completion from a course on ethical treatment of human subjects in research, that you took earlier, you can just attach a copy of that older certificate with your hw, without taking the CITI training course.

(You can use the certificate of completion for several years, in any future human subjects research approval process you may be involved in.)

• Things to read:

– Done already: Groves Ch 1

– This week: Groves, Ch’s 2 and 3

• Clear, careful writing and interpretation of results is an important part of both weekly homeworks and the projects. I always expect neatly typed or neatly handwritten work.

• Always be judicious about including computer output and graphs: show enough that we can clearly see what you are doing, but not so much that we will get lost or bored leafing through your work!

• Mostly homeworks will be submitted on paper in class.

[Exercises begin on next page] Exercises to Turn In (there are 4 excercises):

1. The Health Behavior in School-Age Children Study (HBSC) is an international survey conducted under the auspices of the World Health Organization.

(a) Find the home page for this survey project on the web, and fill out, as completely as possible, a table like Tables 1.1–1.6 of the Groves text, for the the International Report “2001/02 International Report: Young People’s Health in Context”. You may need to use links from the homepage to get the information you need.

(b) Please comment on any items in your table that would make it difficult to understand or believe the results of the survey (e.g. you couldn’t get the information, or the information you got worries you in some way). If you are happy with all the information in your table, just say that.

2. For several years, the American Automobile Association (AAA) has conducted a survey called the Traffic Safety Culture Index. The 2010 report (as well as reports from earlier years) can be found at http://www.aaafoundation.org/reports. Re-do question #1 (both parts!) for the AAA Traffic Safety Culture Index survey.

3. Groves, Ch 2, p. 65, #1. Be sure to give \underline{separate} answers in part (d) for sampling error and for coverage error.

4. Groves, Ch 2, p. 67, #7.